Determine the electrical force of attraction between two balloons with separate charges of +3.5x10-8 C and -2.9x10-8 C when separated a distance of 0.65 m


1
Expert's answer
2021-03-21T11:27:29-0400
"F=\dfrac{kq_1q_2}{r^2},""F=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot3.5\cdot10^{-8}\ C\cdot(-2.9\cdot10^{-8}\ C)}{(0.65\ m)^2}=-2.16\cdot10^{-5}\ N."

The magnitude of the electrical force of attraction between two balloons is "2.16\cdot10^{-5}\ N". Since the two charges are opposite in sign they will attract each other, therefore, the direction of the electrical force is attractive (also, the sign minus indicates that the direction is attractive).